# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v2

@@ D.2 Empirical Transfer Behavior Analysis of the mLSTM
We analyze the transfer behavior of the mLSTM for a single head and a single input sequence of length $T=$512. The inputs are for the queries, keys and values $\bm{q}_{t}$, $\bm{k}_{t}$ and $\bm{v}_{t}$ are sampled from the standard normal distribution $\mathcal{N}(0,1)$. We set the head dimensions to $d_{qk}=$128 and $d_{hv}=$128. As norm layer $\mathrm{NORM}(\bm{x})$ we use the RMS-norm. Changing the norm to layernorm does not alter the results, as for this experiment we set the mean of the inputs to zero. For every plot we measure the gains $G_{\mathrm{before}}$ and $G_{\mathrm{after}}$ (as defined in (26)) for input and forget gate preactivation values in the ranges [-12, 8] and [-5, 12], respectively.
@@
Based on our analysis on the normalization layer after the gated linear RNN operation in Section 4.2, we hypothesize that the normalization layer and especially the norm epsilon $\epsilon$ is integral to the gating mechanism. In this experiment, we probe the effect of the epsilon value on the transfer behavior of the mLSTM. Figure 9(a) and Figure 10(a) show the transfer behavior of mLSTMexp and mLSTMsig for ϵ=italic-$\epsilon=$[1e-2, 1e-6, 1e-8], respectively.
We observe that the epsilon acts in the same way for mLSTMexp and mLSTMsig. Increasing $\epsilon$ causes an offset of the gain in positive y-direction, increasing $\epsilon$ in negative y-direction. We set our default value ϵ=italic-$\epsilon=$1e-6, which yields the best performance in our experiments (see Sec. 5.1).
@@
In contrast, in Figure 10(a) and 10(b) we observe that incorporating a normalizer similar to mLSTMexp (excluding the max state) into mLSTMsig does not alter its transfer behavior.
The other two normalizer variants for mLSTMsig in Figure 10(c) and 10(d) show a clearly different transfer behavior and do not train successfully. Similarly, the variants in Figure 9(b) and 9(c) also fail to train successfully.
@@ E.2 Extended Language Modeling Experiments with mLSTM
An embedding layer, is followed by a stack of blocks and a language model head that produces the output logits (i.e. the values before softmax), which typically consists of a normalization layer and a linear (unembedding) layer. We apply logit soft-capping (Team, 2024), such that the value of the logits stay between $-c$ and $c$ for a specific cap value $c$. We choose $c=$30. The logits are capped with the following function:
@@
Each block consists of two layers, where each layer has skip a connection and a normalization layer before the layer input (i.e. we use the pre-norm block architecture). As normalization layer we use the RMS-norm (Zhang & Sennrich, 2019) with epsilon ϵ=italic-$\epsilon=$1e-6.
@@
For the mLSTM we set the head dimension for the queries and keys to be half of the values, i.e. $d_{qk}=0.5\,d_{hv}$. We use Layernorm (Ba et al., 2016) as $\mathrm{NORM}(\bm{x})$ operation with epsilon ϵ=italic-$\epsilon=$1e-6 in our experiments.  [^1] We confirmed empirically that the type of normalization layer does not affect the performance as well as our qualitative results on transfer behavior and gradient norm variance. Therefore, we generally prefer RMS-norm as it faster. We apply soft-capping from equation (112) on the input and forget gate preactivations, as we found that this improves training stability. For the gate preactivations we set $c=$15.
@@
We train our models with the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=$0.9, $\beta_{2}=$0.95 and ϵ=italic-$\epsilon=$1e-8. We use learning rates and batch sizes as specified in Table 4. We apply a weight decay of 0.1 to all linear layers (including the last linear layer or unembedding) and exclude biases and the token embeddings from weight decay. We clip the gradient norm at 0.5. We use a cosine learning rate scheduler with a linear warmup for the first 750 steps and decay to 0.1 of the peak learning rate, followed by a linear cooldown to 0 for the last 1000 steps. We list the number of training steps for every model size in Table 4. During pre-training we ensure that no information is leaked across document borders by resetting the memory states at the beginning of each new document. We implement this by manually setting the forget gate preactivations to a large negative values at the beginning of each new document.
@@
Effect of Input Gate Bias Initialization (Figure 11 and 12).
In our transfer behavior analysis in Section 4.2 we find that there is a transition from suppressing the signal to passing the signal at negative input gate values of around -8 (see Figure 4). Since we initialize the weights of the gates $\bm{w}_{\{\mathrm{i},\mathrm{f}\}}$ to 0, the biases of the input and forget gates determine the actual position in the x-y plane in the beginning of training. Initially, with input gate biases initialized to 0, we observe a high gradient norm variance, which was more pronounced for mLSTMsig (see Figure 11(a) and 12(a)).
@@
Indeed, as we observe in Figure 11 and 12 initializing the input gate biases to -10 effectively mitigates gradient norm spikes and reduces high gradient norm variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 6), but also improves training stability, if initialized correctly.
@@
Figure 11: Trainable and fixed exponential input gate for bias initializations [0, -2, -5, -10] and norm epsilon ϵ=italic-$\epsilon=$1e-6.
Figure 12: Trainable and fixed sigmoid input gate for bias initializations [0, -2, -5, -10] and norm epsilon ϵ=italic-$\epsilon=$1e-6.
@@
In our empirical transfer behavior analysis of the mLSTM in Section 4.2 and D.2 we find that the transfer behavior depends on the input and forget gate preactivations, as well as the normalization layer epsilon (see Figure 9(a) and 10(a)). Therefore, we perform a grid search over different normalization layer epsilons and input gate bias initializations for the mLSTM with exponential input gate with 160M parameters and 6 heads at context length 4096. We show the results in Figure 13.
@@ E.3 Extended Kernel Benchmark
In Figure 6 and 15 we measure the GPU memory used by the kernels. For this, we use the PyTorch torch.cuda.max_memory_allocated API to measure the peak memory allocated during one kernel iteration. We make sure that the memory statistics are reset after each iteration and that the PyTorch caches are cleared before the start of each benchmark.
@@
We use the standard embedding dimension of 4096 for 7B Transformer models for our benchmark. Since different models and kernels have different default input sizes at this embedding dimension, we adapt the head dimension, number of heads and remaining input dimensions for each kernel accordingly. Following the practice of Shah et al. (2024) we keep the number of tokens constant at 65,536 and vary the sequence length (i.e. $T=[512,1024,2048,4096,8192,16384,32768,65536]$) and batch size accordingly (i.e. $N_{\text{batch}}=65536/T$).
@@
  • Simple GLA (FLA): A simple version of GLA with scalar forget gates per head. This primitive is not published as a new sequence modeling primitive but serves as a reference implementation for kernels for RetNet (Sun et al., 2023) or Mamba 2 (Dao & Gu, 2024) in the FLA library (Yang & Zhang, 2024). Moreover, Simple GLA is similar to mLSTMsig, but has no input gate. Therefore, we find it interesting to add it as baseline. We use the implementation from the official FLA repository, version 0.1.
  • Mamba: Mamba CUDA kernel (Gu & Dao, 2024). Implementation from the official Mamba repository, version 2.2.4.
  • Mamba 2: Mamba 2 Triton kernels (Dao & Gu, 2024). Implementation from the official Mamba repository, version 2.2.4.555See https://github.com/state-spaces/mamba
@@
We show the results of this benchmark for varying sequence length and constant number of tokens in Figure 5. When comparing the forward pass runtime only, we find that Mamba2 and Simple GLA kernels are slightly faster than our mLSTMsig kernels. However, this difference is within 1 ms. In training, when forward and backward pass runtime is measured, our TFLA kernels are faster than FlashAttention 3 for longer sequence lengths and more than two times faster than Mamba 2 kernels for all sequence lengths. Only Simple GLA (FLA) can keep up in training speed with our TFLA mLSTM kernels. Therefore, we compare the runtime and memory usage for a larger head dimension in Figure 15 and find that this comes at the cost of almost 2 times the GPU memory usage compared to our TFLA mLSTM kernels. These memory savings are achieved by leveraging a larger chunk size, enabled through the two levels of sequence parallelism outlined in Section 3.2.
@@ F.1 Exact FLOPs and Memory Operations Count for the mLSTM
Chunkwise-Parallel Formulation (Table 8, 9).
@@
FLOPs mLSTMexp mLSTMsig Recurrent computation of the inter chunk states Gates: $2L+\frac{1}{2}L(L+1)$ $+L(1+F_{\text{exp}}+F_{\text{log}}+F_{\text{sig}})+3+F_{\text{max}}+F_{\text{exp}}$ $2L+\frac{1}{2}L(L+1)+LF_{\text{exp}}+F_{\text{exp}}$ $+2L(F_{\text{log}}+F_{\text{sig}})$ Numerator: $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ Denominator: $2d_{qk}+2Ld_{qk}$ Parallel computation of the intra chunk outputs Cumulative Forget Gates: $\frac{1}{2}L(L+1)+L(F_{\text{log}}+F_{\text{sig}})$ $\frac{1}{2}L(L+1)+2L(F_{\text{log}}+F_{\text{sig}})$ Gate Matrix: $F_{\text{causal}}\times\left(L^{2}(3+F_{\text{exp}}+F_{\text{max}})+L(1+F_{\text{max}})\right)$ $F_{\text{causal}}\times\left(L^{2}(2+F_{\text{exp}})\right)$ Intra Outputs: $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ Parallel computation of the inter chunk outputs Inter Outputs: $2Ld_{qk}d_{hv}+3Ld_{qk}$ $2Ld_{qk}d_{hv}+Ld_{qk}$ Combination of inter and intra chunk outputs Output Combination: $2Ld_{hv}+L(1+F_{\text{max}}+F_{\text{abs}}+F_{\text{exp}})$ $Ld_{hv}$
@@
Load: $L(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2L\times\text{bytes}_{if}$ $+(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cnm}$ $L(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2L\times\text{bytes}_{if}$ $+d_{qk}d_{hv}\times\text{bytes}_{Cnm}$
@@
Total: $4L\times\text{bytes}_{if}$ $+3L\left(d_{hv}+d_{qk}\right)\times\text{bytes}_{qkv}$ $+2\left(L+d_{hv}d_{qk}+d_{qk}+1\right)\times\text{bytes}_{Cmn}$ $4L\times\text{bytes}_{if}$ $+3L\left(d_{hv}+d_{qk}\right)\times\text{bytes}_{qkv}$ $+2d_{hv}d_{qk}\times\text{bytes}_{Cmn}$
Fully Parallel Formulation (Table 10, 11).
@@
FLOPs mLSTMexp mLSTMsig
Cumulative Forget Gates: $\frac{1}{2}T(T+1)+T(F_{\text{log}}+F_{\text{sig}})$ $\frac{1}{2}T(T+1)+2T(F_{\text{log}}+F_{\text{sig}})$
Gate Matrix: $T^{2}(3+F_{\text{exp}}+F_{\text{max}}+F_{\text{mask}})$ $T^{2}(3+F_{\text{exp}}+F_{\text{max}}+F_{\text{mask}})$
Normalization: $F_{\text{causal}}\times\left(T^{2}(3+F_{\text{abs}})+T(F_{\text{exp}}+F_{\text{max}})\right)$
@@
Bytes mLSTMexp mLSTMsig
Load: $T(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2T\times\text{bytes}_{if}$ $T(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2T\times\text{bytes}_{if}$
Store: $Td_{hv}\times\text{bytes}_{qkv}+2T\times\text{bytes}_{Cmn}$ $Td_{hv}\times\text{bytes}_{qkv}$
Total: $2T(\text{bytes}_{if}+(d_{hv}+d_{qk})\times\text{bytes}_{qkv}+\text{bytes}_{Cmn})$ $2T(\text{bytes}_{if}+(d_{hv}+d_{qk})\times\text{bytes}_{qkv})$
Recurrent Formulation (Table 12, 13).
@@
Bytes mLSTMexp mLSTMsig
Load: $(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2\times\text{bytes}_{if}$ $+(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cmn}$ $(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2\times\text{bytes}_{if}$ $+d_{qk}d_{hv}\times\text{bytes}_{Cmn}$
Store: $d_{hv}\times\text{bytes}_{qkv}+(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cmn}$ $d_{hv}\times\text{bytes}_{qkv}+d_{qk}d_{hv}\times\text{bytes}_{Cmn}$
Total: $2\times\text{bytes}_{if}+2(d_{hv}+d_{qk})\times\text{bytes}_{qkv}$ $+2d_{hv}d_{qk}\times\text{bytes}_{Cmn}$ $2\times\text{bytes}_{if}+2(d_{hv}+d_{qk})\times\text{bytes}_{qkv}$ $+2(d_{hv}d_{qk}+d_{qk}+1)\times\text{bytes}_{Cmn}$

### v3

@@ D.2 Empirical Transfer Behavior Analysis of the mLSTM
We analyze the transfer behavior of the mLSTM for a single head and a single input sequence of length T=T=512. The inputs are for the queries, keys and values $\bm{q}_{t}$, $\bm{k}_{t}$ and $\bm{v}_{t}$ are sampled from the standard normal distribution $\mathcal{N}(0,1)$. We set the head dimensions to $d_{qk}=$128 and $d_{hv}=$128. As norm layer $\mathrm{NORM}(\bm{x})$ we use the RMS-norm. Changing the norm to layernorm does not alter the results, as for this experiment we set the mean of the inputs to zero. For every plot we measure the gains $G_{\mathrm{before}}$ and $G_{\mathrm{after}}$ (as defined in (16)) for input and forget gate preactivation values in the ranges [-12, 8] and [-5, 12], respectively.
@@
Based on our analysis on the normalization layer after the gated linear RNN operation in Section 4.2, we hypothesize that the normalization layer and especially the norm epsilon $\epsilon$ is integral to the gating mechanism. In this experiment, we probe the effect of the epsilon value on the transfer behavior of the mLSTM. Figure 10(a) and Figure 11(a) show the transfer behavior of mLSTMexp and mLSTMsig for $\epsilon=$[1e-2, 1e-6, 1e-8], respectively.
We observe that the epsilon acts in the same way for mLSTMexp and mLSTMsig. Increasing $\epsilon$ causes an offset of the gain in positive y-direction, increasing $\epsilon$ in negative y-direction. We set our default value $\epsilon=$1e-6, which yields the best performance in our experiments (see Sec. 5.1).
@@
In contrast, in Figure 11(a) and 11(b) we observe that incorporating a normalizer similar to mLSTMexp (excluding the max state) into mLSTMsig does not alter its transfer behavior.
The other two normalizer variants for mLSTMsig in Figure 11(c) and 11(d) show a clearly different transfer behavior and do not train successfully. Similarly, the variants in Figure 10(b) and 10(c) also fail to train successfully.
@@ E.2 Extended Language Modeling Experiments with mLSTM
An embedding layer, is followed by a stack of blocks and a language model head that produces the output logits (i.e. the values before softmax), which typically consists of a normalization layer and a linear (unembedding) layer. We apply logit soft-capping (Team, 2024), such that the value of the logits stay between c-c and $c$ for a specific cap value $c$. We choose c=c=30. The logits are capped with the following function:
@@
Each block consists of two layers, where each layer has skip a connection and a normalization layer before the layer input (i.e. we use the pre-norm block architecture). As normalization layer we use the RMS-norm (Zhang & Sennrich, 2019) with epsilon $\epsilon=$1e-6.
@@
For the mLSTM we set the head dimension for the queries and keys to be half of the values, i.e. $d_{qk}=0.5\,d_{hv}$. We use Layernorm (Ba et al., 2016) as $\mathrm{NORM}(\bm{x})$ operation with epsilon $\epsilon=$1e-6 in our experiments.  [^1] We confirmed empirically that the type of normalization layer does not affect the performance as well as our qualitative results on transfer behavior and gradient norm variance. Therefore, we generally prefer RMS-norm as it faster. We apply soft-capping from equation (102) on the input and forget gate preactivations, as we found that this improves training stability. For the gate preactivations we set c=c=15.
@@
We train our models with the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=$0.9, $\beta_{2}=$0.95 and $\epsilon=$1e-8. We use learning rates and batch sizes as specified in Table 4. We apply a weight decay of 0.1 to all linear layers (including the last linear layer or unembedding) and exclude biases and the token embeddings from weight decay. We clip the gradient norm at 0.5. We use a cosine learning rate scheduler with a linear warmup for the first 750 steps and decay to 0.1 of the peak learning rate, followed by a linear cooldown to 0 for the last 1000 steps. We list the number of training steps for every model size in Table 4. During pre-training we ensure that no information is leaked across document borders by resetting the memory states at the beginning of each new document. We implement this by manually setting the forget gate preactivations to a large negative values at the beginning of each new document.
@@
Table 6: Training Step Time in seconds at context length 8192. EXP and SIG denote mLSTMexp and mLSTMsig. LIMIT and XL correspond to limit_chunk and xl_chunk kernels. All runs use the same training setup with global batch size of 256 sharded across 32 GPUs corresponding to a local batch size of 8.
Size Batch Size Global / Local Heads Llama EXP LIMIT EXP XL SIG XL 1.4B 256 / 8 4 1.77 1.73 1.67 8 1.67 1.66 1.64 16 2.39 1.60 1.61 1.62
@@
Effect of Input Gate Bias Initialization (Figure 12 and 13).
In our transfer behavior analysis in Section 4.2 we find that there is a transition from suppressing the signal to passing the signal at negative input gate values of around -8 (see Figure 4). Since we initialize the weights of the gates $\bm{w}_{\{\mathrm{i},\mathrm{f}\}}$ to 0, the biases of the input and forget gates determine the actual position in the x-y plane in the beginning of training. Initially, with input gate biases initialized to 0, we observe a high gradient norm variance, which was more pronounced for mLSTMsig (see Figure 12(a) and 13(a)).
@@
Indeed, as we observe in Figure 12 and 13 initializing the input gate biases to -10 effectively mitigates gradient norm spikes and reduces high gradient norm variance during training for both mLSTMexp and mLSTMsig. We therefore conclude that the additional input gate not only improves performance (see Table 7), but also improves training stability, if initialized correctly.
@@
Figure 12: Trainable and fixed exponential input gate for bias initializations [0, -2, -5, -10] and norm epsilon $\epsilon=$1e-6.
Figure 13: Trainable and fixed sigmoid input gate for bias initializations [0, -2, -5, -10] and norm epsilon $\epsilon=$1e-6.
@@
In our empirical transfer behavior analysis of the mLSTM in Section 4.2 and D.2 we find that the transfer behavior depends on the input and forget gate preactivations, as well as the normalization layer epsilon (see Figure 10(a) and 11(a)). Therefore, we perform a grid search over different normalization layer epsilons and input gate bias initializations for the mLSTM with exponential input gate with 160M parameters and 6 heads at context length 4096. We show the results in Figure 14.
@@ E.3 Extended Kernel Benchmark
In Figure 6 and 16 we measure the GPU memory used by the kernels. For this, we use the PyTorch torch.cuda.max_memory_allocated API to measure the peak memory allocated during one kernel iteration. We make sure that the memory statistics are reset after each iteration and that the PyTorch caches are cleared before the start of each benchmark.
@@
We use the standard embedding dimension of 4096 for 7B Transformer models for our benchmark. Since different models and kernels have different default input sizes at this embedding dimension, we adapt the head dimension, number of heads and remaining input dimensions for each kernel accordingly. Following the practice of Shah et al. (2024) we keep the number of tokens constant at 65,536 and vary the sequence length (i.e. T=[512,1024,2048,4096,8192,16384,32768,65536]T=[512,1024,2048,4096,8192,16384,32768,65536]) and batch size accordingly (i.e. $N_{\text{batch}}=65536/T$).
@@
  • Simple GLA (FLA): A simple version of GLA with scalar forget gates per head. This primitive is not published as a new sequence modeling primitive but serves as a reference implementation for kernels for RetNet (Sun et al., 2023) or Mamba 2 (Dao & Gu, 2024) in the FLA library Yang & Zhang (2024). Moreover, Simple GLA is similar to mLSTMsig, but has no input gate. Therefore, we find it interesting to add it as baseline. We use the implementation from the official FLA repository, version 0.1.
  • Mamba: Mamba CUDA kernel Gu & Dao (2024). Implementation from the official Mamba repository, version 2.2.4.
  • Mamba 2: Mamba 2 Triton kernels Dao & Gu (2024). Implementation from the official Mamba repository, version 2.2.4.555See https://github.com/state-spaces/mamba
@@
We show the results of this benchmark for varying sequence length and constant number of tokens in Figure 5. When comparing the forward pass runtime only, we find that Mamba2 and Simple GLA kernels are slightly faster than our mLSTMsig kernels. However, this difference is within 1 ms. In training, when forward and backward pass runtime is measured, our TFLA kernels are faster than FlashAttention 3 for longer sequence lengths and more than two times faster than Mamba 2 kernels for all sequence lengths. Only Simple GLA (FLA) can keep up in training speed with our TFLA mLSTM kernels. Therefore, we compare the runtime and memory usage for a larger head dimension in Figure 16 and find that this comes at the cost of almost 2 times the GPU memory usage compared to our TFLA mLSTM kernels. These memory savings are achieved by leveraging a larger chunk size, enabled through the two levels of sequence parallelism outlined in Section 3.
@@ F.1 Exact FLOPs and Memory Operations Count for the mLSTM
Chunkwise-Parallel Formulation (Table 9, 10).
@@
FLOPs mLSTMexp mLSTMsig Recurrent computation of the inter chunk states Gates: $2L+\frac{1}{2}L(L+1)$ $L(1+F_{\text{exp}}+F_{\text{log}}+F_{\text{sig}})+3+F_{\text{max}}+F_{\text{exp}}$ $2L+\frac{1}{2}L(L+1)+LF_{\text{exp}}+F_{\text{exp}}$ $2L(F_{\text{log}}+F_{\text{sig}})$ Numerator: $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ Denominator: $2d_{qk}+2Ld_{qk}$ Parallel computation of the intra chunk outputs Cumulative Forget Gates: $\frac{1}{2}L(L+1)+L(F_{\text{log}}+F_{\text{sig}})$ $\frac{1}{2}L(L+1)+2L(F_{\text{log}}+F_{\text{sig}})$ Gate Matrix: $F_{\text{causal}}\times\left(L^{2}(3+F_{\text{exp}}+F_{\text{max}})+L(1+F_{\text{max}})\right)$ $F_{\text{causal}}\times\left(L^{2}(2+F_{\text{exp}})\right)$ Intra Outputs: $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ Parallel computation of the inter chunk outputs Inter Outputs: $2Ld_{qk}d_{hv}+3Ld_{qk}$ $2Ld_{qk}d_{hv}+Ld_{qk}$ Combination of inter and intra chunk outputs Output Combination: $2Ld_{hv}+L(1+F_{\text{max}}+F_{\text{abs}}+F_{\text{exp}})$ $Ld_{hv}$
@@
Load: $L(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2L\times\text{bytes}_{if}$ $(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cnm}$ $L(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2L\times\text{bytes}_{if}$ $d_{qk}d_{hv}\times\text{bytes}_{Cnm}$
@@
Total: $4L\times\text{bytes}_{if}$ $3L\left(d_{hv}+d_{qk}\right)\times\text{bytes}_{qkv} +$$2\left(L+d_{hv}d_{qk}+d_{qk}+1\right)\times\text{bytes}_{Cmn}$ $4L\times\text{bytes}_{if}$ $3L\left(d_{hv}+d_{qk}\right)\times\text{bytes}_{qkv} +$$2d_{hv}d_{qk}\times\text{bytes}_{Cmn}$
Fully Parallel Formulation (Table 11, 12).
@@
FLOPs mLSTMexp mLSTMsig Cumulative Forget Gates: $\frac{1}{2}T(T+1)+T(F_{\text{log}}+F_{\text{sig}})$ $\frac{1}{2}T(T+1)+2T(F_{\text{log}}+F_{\text{sig}})$ Gate Matrix: $T^{2}(3+F_{\text{exp}}+F_{\text{max}}+F_{\text{mask}})$ $T^{2}(3+F_{\text{exp}}+F_{\text{max}}+F_{\text{mask}})$ Attention Logits: $F_{\text{causal}}\times\left(2T^{2}d_{qk}+2T^{2}\right)$ $F_{\text{causal}}\times\left(2T^{2}d_{qk}+2T^{2}\right)$ Normalization: $F_{\text{causal}}\times\left(T^{2}(3+F_{\text{abs}})+T(F_{\text{exp}}+F_{\text{max}})\right)$ Outputs: $F_{\text{causal}}\times 2T^{2}d_{hv}$ $F_{\text{causal}}\times 2T^{2}d_{hv}$
@@
Bytes mLSTMexp mLSTMsig Load: $T(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2T\times\text{bytes}_{if}$ $T(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2T\times\text{bytes}_{if}$ Store: $Td_{hv}\times\text{bytes}_{qkv}+2T\times\text{bytes}_{Cmn}$ $Td_{hv}\times\text{bytes}_{qkv}$ Total: $2T(\text{bytes}_{if}+(d_{hv}+d_{qk})\times\text{bytes}_{qkv}+\text{bytes}_{Cmn})$ $2T(\text{bytes}_{if}+(d_{hv}+d_{qk})\times\text{bytes}_{qkv})$
Recurrent Formulation (Table 13, 14).
@@
Bytes mLSTMexp mLSTMsig Load: $(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2\times\text{bytes}_{if}$ $(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cmn}$ $(2d_{qk}+d_{hv})\times\text{bytes}_{qkv}+2\times\text{bytes}_{if}$ $d_{qk}d_{hv}\times\text{bytes}_{Cmn}$ Store: $d_{hv}\times\text{bytes}_{qkv}+(d_{qk}d_{hv}+d_{qk}+1)\times\text{bytes}_{Cmn}$ $d_{hv}\times\text{bytes}_{qkv}+d_{qk}d_{hv}\times\text{bytes}_{Cmn}$ Total: $2\times\text{bytes}_{if}+2(d_{hv}+d_{qk})\times\text{bytes}_{qkv}$ $2d_{hv}d_{qk}\times\text{bytes}_{Cmn}$ $2\times\text{bytes}_{if}+2(d_{hv}+d_{qk})\times\text{bytes}_{qkv}$ $2(d_{hv}d_{qk}+d_{qk}+1)\times\text{bytes}_{Cmn}$